Given a random variable x distributed as N (5, 64; x), determine the following probability

P(|x−7|≤1)

For the random variable in the previous problem N (5, 64; x), determine the values Zi such that

P( x≥ Z2)=0.11
Answer #1

PL lx-7|c1) P (-(x-7)く12 х-7) ,. = = 0.6962-00013 z =-1 . 227